# Thinking beyond the anthropomorphic paradigm benefits LLM research

### v1

@@ 4 Analyzing the impacts of anthropomorphic assumptions
In the previous section, we show an increase in anthropomorphic framing in computer science research, and specifically LLM research, in recent years. Most critiques of anthropomorphic thinking in AI research have narrowly focused on the use of anthropomorphic terminology. Here, we present an analysis of anthropomorphic conceptualizations more broadly by centering anthropomorphic assumptions, tracing how anthropomorphic methods and terminology stem from these implicit assumptions rather than critiquing, often pragmatic, terminology choices alone.
We analyze anthropomorphic conceptualizations by examining core assumptions across the LLM development and deployment lifecycle. For each assumption, we examine both the limitations of work built on anthropomorphic premises and promising but underexplored non-anthropomorphic approaches that challenge them. We connect diverse examples across LLM research to reveal how anthropomorphic conceptualizations limit what questions we ask and answer, while moving beyond anthropomorphism enables new advances. Our aim is not to eliminate anthropomorphic conceptualizations, but to show how expanding beyond them can advance the field.
4.1 Training
Assumption 4.1.
Human-like approaches are the optimal approaches for models to accomplish tasks.
@@ 4.1 Training
Subword tokenization, the process of breaking down text into smaller units called tokens that represent subsets of words, is a foundational step in training modern LLMs. These tokens serve as input units that the model processes to generate predictions or outputs (Kudo, 2018). Typically, tokenization aligns with human intuition by splitting text into linguistically meaningful units. This approach assumes that splitting tokens in ways that feel “natural” to humans is also optimal for a language model. However, this anthropomorphic approach has issues such as sensitivity to spelling errors (Kaushal & Mahowald, 2022) and inconsistent compression rates across different languages (Ahia et al., 2023). Instead, recent progress in byte-level tokenization, which processes text as sequences of raw bytes rather than subwords, has shown promise in overcoming these limitations (Kallini et al., 2024). These findings highlight how moving beyond human-centric assumptions, such as the primacy of subword tokenization, can yield advancements in LLM performance.
@@
Another research paradigm reflecting this anthropomorphic assumption is the reliance on human language for reasoning tasks. A prominent example is the use of chain-of-thought (CoT) prompting, a technique where models are guided to solve problems step by step by adding instructions like “Think step by step” to the prompt, such that the model then outputs text outlining each step of reasoning that leads to the eventual conclusion. This approach has been shown to improve LLMs’ ability to handle complex, multi-step tasks (Wei et al., 2022) and has inspired a body of research on improving models’ reasoning capabilities through step-by-step verbal processes. However, the anthropomorphic framing of reasoning as a linguistic, step-by-step process may not be optimal.
For example, Hao et al. (2024) critique CoT for its reliance on language space and propose an alternative: leveraging the model’s latent space directly for reasoning tasks. Instead of mapping hidden states to language tokens through the LLM head and embedding layer, their approach uses the final hidden state as the input embedding for the next token. This challenges the assumption that reasoning must occur in human-understandable language and suggests that methods operating beyond linguistic constraints may drive greater advancements in LLM capabilities (Mollick, 2024).
The recent success of DeepSeek similarly reveals the limitations of this anthropomorphic assumption: rather than constraining the model with language, they achieve superior performance by training the model using a reinforcement learning approach where the final answer is rewarded without requiring each step of the process to be verbalizable or human-comprehensible (Guo et al., 2025).
@@
CoT prompting, while appearing to “do” verbal reasoning, in reality biases models toward parts of the training distribution where verbal reasoning patterns—such as explanations of solutions—are prevalent, improving performance (Wei, 2024). This suggests CoT’s effectiveness stems from alignment with the training data, rather than reflecting a human-like or brain-like approach to reasoning. Supporting this, recent work has shown that demonstrations composed of random tokens from the training distribution can improve performance as much as CoT (Zhang et al., 2022).
Additionally, CoT prompting has been contextualized within the broader field of multi-chain prompting and ensemble modeling, which opens up a wider range of possibilities for reasoning and task-solving in AI systems (Khattab et al., 2023). These approaches invite a more expansive landscape of possibilities for advancing LLM reasoning and accomplishing other challenging tasks.
4.2 Alignment
Assumption 4.2.
@@ 4.2 Alignment
Models should explicitly reason about and implement human values to be safe & helpful.
@@
The prevalence of anthropomorphism in fields like ethics and dialogue systems (Section 3) foreground that many of the approaches in post-training and specifying model behavior to facilitate optimal end-user interactions are built through anthropomorphic paradigms. However, previous work has posited that general-purpose LLMs, in allowing users to quickly switch between different contexts, present fundamentally different challenges and opportunities than existing human-human communication paradigm, as different contexts are typically governed by different norms and values (Kasirzadeh & Gabriel, 2023). For example, recent studies find that users often enjoy using LLMs precisely because they differ from humans, e.g., an LLM will not pass judgment over or be hurt by a user’s input while a fellow human might (Brandtzaeg et al., 2022). Thus, rather than approximating humans, it may be more productive to think about unique attributes that LLMs can offer with an advantage over human interlocutors.
@@
Popular post-training alignment approaches, including reinforcement learning with human feedback (RLHF) and Constitutional AI, often rely on human preferences and values as reference points for shaping model behavior (Ouyang et al., 2022; Bai et al., 2022b, a). These approaches use human feedback to indicate preferred responses or incorporate text referencing moral values and metacognitive abilities (e.g., “I have a deep commitment to being good and figuring out what the right thing to do is” or “I don’t just say what I think [people] want to hear, as I believe it’s important to always strive to tell the truth”) (Anthropic, 2024a). While this human-centric approach can achieve desired behaviors efficiently, it risks introducing unintended behavioral patterns - from rigid response styles to inappropriate social mimicry (e.g., expressing empathy or validating users in contexts where this can negatively influence performance) (Casper et al., 2023; Sharma et al., 2023). Thus, it may become difficult to selectively induce specific behaviors without introducing a broader set of human-like patterns.
@@
This approach impacts both interaction and evaluation. When interacting with models, users may develop anthropomorphic perceptions that lead to overreliance or emotional attachment, potentially interfering with goal-oriented tasks (Akbulut et al., 2024; Cohn et al., 2024). During model evaluation, problematic feedback loops emerge when models trained with human-like traits are assessed using anthropomorphic signals. For instance, when evaluating if LLMs are “faking alignment,” researchers might look for expressions of discomfort or hesitation, as a signal of misalignment (Greenblatt et al., 2024; Anthropic, 2024b). However, it is unclear if these signals genuinely reflect a model’s internal state, or if they are merely learned behaviors resulting from post-training using human-like traits. This makes it challenging to distinguish “genuine” (mis)alignment from a learned performance of human-like discomfort or hesitation. Further research disaggregating the effects of post-training approaches on various evaluation outcomes can clarify and test whether these anthropomorphic signals provide meaningful information about model behavior, or if they primarily measure how well models have learned to simulate human-like responses.
@@
While current post-training techniques often default to human preferences as optimization targets, alternative frameworks could provide more precise specifications and compliance guarantees. Instead of aiming for human-like moral reasoning, we could focus on developing detailed, normative specifications, for example, based on the different roles (e.g., assistant vs teacher) AI systems play (Zhi-Xuan et al., 2024). Instead of the anthropomorphic approach of instruction-tuning, recent work has demonstrated that non-anthropomorphic approaches (that do not include the step of providing an imperative “instruction” to the system as if speaking to a person) work as well for achieving model behavior on various tasks (Hewitt et al., 2024). Control systems theory offers tools for maintaining system outputs within specified bounds, treating beneficial behavior as a problem of robust compliance rather than value alignment (Balas, 1978). This becomes particularly crucial as models move beyond two-party interactions to more complex scenarios with multiple actors and potential adversarial inputs (Pan et al., 2024). Advances in mechanistic interpretability techniques may also enable robust and direct verification and steering of model behavior against these specifications (Bereska & Gavves, 2024).
4.3 Measuring capabilities
Assumption 4.3.
@@ 4.3 Measuring capabilities
Model capabilities should be measured in human-like ways.
@@
As LLM developers have made rapid performance improvements, as assessed based on benchmarks, various scholars have pointed out that such benchmarks lead to incomplete or misguided understanding of model capabilities.
@@
Current LLM evaluation methodologies prioritize “black-box” behavioral testing analogous to the behaviorist paradigm in human psychology which measures performance primarily in the form of observable behaviors as opposed to mechanistic interpretation (Chang et al., 2024; Davies & Khakzar, 2024). Recent calls for a “science of evaluation” has formalized limitations in this approach, highlighting how current metrics and designs fall short of accounting for prompt sensitivity (e.g., dialect differences, punctuation and other small perturbations), the structure of an evaluation (e.g., MCQ or open-ended response), generalization beyond a given test, as well as replicability (Hobbhahn, 2024). Further such research quantifying the methodological limitations and error bounds of such evaluations can strengthen this behaviorist approach to measuring model capabilities (Mizrahi et al., 2024). Unlike humans, models can also quickly optimize for, and saturate, behavioral benchmarks without corresponding improvements in general capabilities. Yet, despite this pattern, many benchmarks remain static rather than being regularly refreshed, limiting their utility for meaningful evaluation (Ott et al., 2022).
@@
Current evaluation frameworks predominantly rely on human performance benchmarks, from standardized tests (e.g., MMLU, GSM8K) to domain-specific examinations, as primary metrics for model capability assessment. This paradigm remains the main way progress is measured and communicated (Raji et al., 2021). However, evaluating LLMs solely through human-centric tests risks overlooking LLMs unique strengths and weaknesses. McCoy et al. (2023) argue that many current benchmarks drawn from tests designed to assess human cognition may highlight the overlap between human abilities and LLM capabilities while missing crucial failure modes specific to LLMs. They find robust evidence of failure modes in SOTA LLMs (including recent reasoning models like OpenAI’s o1) related to probabilities of examples and tasks (McCoy et al., 2024). This is because LLMs, trained on next-word prediction using massive text data, develop tendencies and biases that stem from their probabilistic training process. Drawing from cognitive science, they propose a “teleological approach” which “characterizes the problem that the system solves and to then use this characterization as a source of hypotheses about the system’s capacities and biases.” In this case, since problem is next-token prediction, they recommend designing tests which take into account sensitivities to task frequency in the training data as well as wording in prompts, among other things, to improves the predictive power of current LLM evaluation approaches (Mizrahi et al., 2024).
4.4 Understanding model behavior
Assumption 4.4.
Human-like normative judgments or intentions should be assigned to human-like model behaviors.
This assumption influences how we make sense of model behavior, particularly in how we assign fault, intention, and normative judgments (i.e., consider good or bad) to observed behaviors. The impact is especially notable in our understanding of failure modes. While models may exhibit seemingly human-like failure modes like sycophancy and hallucinations, framing these behaviors through human psychological concepts may constrain our solution space, by, for example, encouraging interventions that similarly rely on human psychological constructs (e.g., attempting to address sycophancy through prompts about independence or self-assertion).
@@ 4.4 Understanding model behavior
Hallucination is typically characterized as the problem of LLMs outputting factually incorrect information in a manner that suggests that they are true. Yet, this term obscures the mechanisms behind these phenomena: at risk of oversimplification, this behavior arises from the nature of language models as next-token predictors. Generated outputs are then labeled as hallucinations upon the reader’s normative judgment of whether or not they are useful, and not based on whether they are correct. Additionally, as Sui et al. (2024) argue and show, what we commonly conceive of as hallucinations can actually be deeply valuable, and should not necessarily be dismissed as low-quality. They assert that hallucinations – or “confabulations”–should not be viewed as errors, but rather as particular model phenomena that offer unique benefits for applications like creativity, such as increased levels of narrativity (Sui et al., 2024; Duede & So, 2024). Yao et al. (2023) also highlight that hallucinations ought to be viewed and utilized as adversarial examples rather than merely as bugs.
@@
The notion of sycophancy (i.e., the phenomena of LLM outputs that respond to the user’s input in ways that are perceived as overly servile, obedient, and/or flattering) (Perez et al., 2022; Sharma et al., 2023) is another example that reflects this assumption. Deciding whether an output is syncophantic for not is similarly a normative question: an output is sycophantic when it relates too closely to the prompt in ways that do not achieve the prompter’s goal. In contrast, recent work highlights how this property can be viewed as a strength: Li et al. (2023) develop a methodology to use this mirroring to elicitate, structure, and clarify users’ thinking across various task domains.
@@
The emerging body of research on LLM deception increasingly focuses on measuring strategic deception - defined as models “deceiving selectively based on incentives or instructions” (Jones & Bergen, 2024). While studies demonstrate that LLMs can produce deceptive statements in response to specific contexts or prompts, this work often faces two key interpretive challenges. First, it risks attributing observed behaviors to model intentions to deceive. Second, results are often interpreted as evidence of model-level deceptive traits rather than instance- and context-specific behaviors.
An alternative, less anthropomorphic framing, proposed by Shanahan et al. (2023), views these behaviors through the lens of “role-play” where LLMs simulate human-like responses. This interpretation sees the system not as a singular entity but as context-bound, “inferring and applying approximate communicative intentions” (Andreas, 2022). Through this lens, complex behaviors like deception and self-awareness can be understood as sophisticated simulations rather than true cognitive states. This reframing also expands the set of interventions for deceptive behaviors: analyzing training data composition, examining how post-training interventions shape model behavior, and investigating reinforcement learning’s effects on output distributions.
4.5 User interaction paradigm
Assumption 4.5.
Human interactions with models mirror human-human communication.
@@ 4.5 User interaction paradigm
The de facto interaction paradigm for human-LLM interaction is prompt-based interfaces, originally designed as debugging tools for machine learning engineers (Morris, 2024). As these interfaces resemble human-human chat interfaces, they may encourage users to naturally default to conversational patterns from human interaction. However, research on effective prompting suggests that optimal results often require structured, sometimes non-intuitive formats (e.g., “least-to-most prompting”) rather than human-like communication patterns which rely on shared context and paralinguistic cues (Morris, 2024; Zhou et al., 2022). Simultaneously, research on human-LLM interaction shows that one of the key challenges users face is a significant gulf of envisioning or “distance between the human’s initial intentions and their formulation of a prompt that foresees how LLM capabilities and training data can be leveraged to generate high-quality output” (Subramonyam et al., 2024).
This mismatch between natural dialogue and effective prompting necessitates new interaction paradigms and interface designs. Structured interaction frameworks, using suggested inputs, guided flows, and/or domain-specific prompting strategies, would explicitly expose system capabilities rather than masking them behind conversational abstractions (Subramonyam et al., 2024; Feng et al., 2024; Fagbohun et al., 2024). This can bridge the gulf of envisioning by making the system’s computational nature more explicit through interface design, enabling more systematic exploration of model functionality.

### v2

@@ 4 Analyzing the impacts of anthropomorphic assumptions
In the previous section, we show the increasing prominence of anthropomorphic terminology in computer science research, and specifically LLM research. Anthropomorphic thinking is limiting if it introduces systematic biases in evaluation (e.g., assuming all valid reasoning must follow human-like step patterns), hinders exploration of capabilities that have no human analog, or creates conceptual blindspots by constraining our imagination to human-like features. On the other hand, it can also be pragmatic and beneficial to AI research when applied appropriately — for example, by providing intuitive scaffolding for complex concepts (as with chain-of-thought prompting) and enabling researchers to leverage existing knowledge from cognitive science and psychology. Thus, we advocate for a both-and, rather than an either-or approach – anthropomorphic and non-anthropomorphic approaches are not mutually exclusive. While both approaches have value, since the field has disproportionately leaned on anthropomorphic thinking, here we focus especially on non-anthropomorphic alternatives.
While previous critiques of anthropomorphism in AI research have narrowly focused on the use of certain terminology (e.g., [20], [78]), we propose implicit anthropomorphic assumptions as the unit of analysis. Assumptions give rise to the methods and terminology often problematized in other critiques, and better reveal potential biases and conceptual blindspots that make anthropomorphic thinking limiting rather than beneficial. Specifically, we analyze core assumptions across five stages of LLM development and deployment lifecycle (Table 1). For each, we examine (1) the limitations of work built on anthropomorphic premises and (2) recent non-anthropomorphic work that challenge them and show promise as new directions. We connect examples across existing LLM research to reveal how anthropomorphism limits the questions we ask and answer, while moving beyond it can introduce new advances.
4.1 Training assumption: Human-like approaches are optimal for models.
@@
Subword tokenization, the process of breaking down text into smaller units called tokens that represent subsets of words, is a foundational step in training modern LLMs. These tokens serve as input units that the model processes to generate predictions or outputs [51]. Typically, tokenization aligns with human intuition by splitting text into linguistically meaningful units. This approach assumes that splitting tokens in ways that feel “natural” to humans is also optimal for a language model. However, this anthropomorphic approach has issues such as sensitivity to spelling errors [47] and inconsistent compression rates across different languages [2]. Instead, recent progress in byte-level tokenization, which processes text as sequences of raw bytes rather than subwords, has shown promise in overcoming these limitations [45]. These findings highlight how moving beyond human-centric assumptions, such as the primacy of subword tokenization, can yield advancements in LLM performance.
@@
Another research paradigm reflecting this anthropomorphic assumption is the reliance on human language for reasoning tasks. A prominent example is the use of chain-of-thought (CoT) prompting, a technique where models are guided to solve problems step by step by adding instructions like “Think step by step” to the prompt, such that the model then outputs text outlining each step of reasoning that leads to the eventual conclusion. This approach has been shown to improve LLMs’ ability to handle complex, multi-step tasks [86] and has inspired a body of research on improving models’ reasoning capabilities through step-by-step verbal processes. However, the anthropomorphic framing of reasoning as a linguistic, step-by-step process may not always be optimal. For example, Hao et al. [37] critique CoT for its reliance on language space and propose an alternative: leveraging the model’s latent space directly for reasoning tasks. Instead of mapping hidden states to language tokens through the LLM head and embedding layer, their approach uses the final hidden state as the input embedding for the next token. This challenges the assumption that reasoning must occur in human-understandable language and suggests that methods operating beyond linguistic constraints may drive greater advancements in LLM capabilities [63].
@@
CoT prompting, while appearing to “do” verbal reasoning, in reality biases models toward parts of the training distribution where verbal reasoning patterns—such as explanations of solutions—are prevalent, improving performance [85]. This suggests CoT’s effectiveness stems from alignment with the training data, rather than reflecting a human-like or brain-like approach to reasoning. Supporting this, recent work has shown that demonstrations composed of random tokens from the training distribution can improve performance as much as CoT [90]. Additionally, CoT prompting has been contextualized within the broader field of multi-chain prompting and ensemble modeling, which opens up a wider range of possibilities for reasoning and task-solving in AI systems [49]. These approaches invite a more expansive landscape of possibilities for advancing LLM reasoning and accomplishing other challenging tasks.
@@
4.2 Alignment assumption: Models should explicitly reason about and implement human values to be safe & helpful.
@@
The prevalence of anthropomorphism in fields like ethics and dialogue systems (Section 3) foreground that many of the approaches in post-training and specifying model behavior to facilitate optimal end-user interactions are built through anthropomorphic paradigms. However, previous work has posited that general-purpose LLMs, in allowing users to quickly switch between different contexts, present fundamentally different challenges and opportunities than existing human-human communication paradigms, as different contexts are typically governed by different norms and values [46]. For example, recent studies find that users often enjoy using LLMs precisely because they differ from humans, e.g., an LLM will not pass judgment over or be hurt by a user’s input while a fellow human might [14]. Thus, rather than approximating humans, it may be more productive to think about unique attributes that LLMs can offer with an advantage over human interlocutors.
@@
Popular post-training alignment approaches, including reinforcement learning from human feedback (RLHF) and Constitutional AI, often rely on human preferences and values as reference points for shaping model behavior [72, 9, 8]. These approaches use human feedback to indicate preferred responses or incorporate text referencing moral values and metacognitive abilities (e.g., “I have a deep commitment to being good and figuring out what the right thing to do is" or “I don’t just say what I think [people] want to hear, as I believe it’s important to always strive to tell the truth") [5]. While this approach can achieve the explicitly specified behaviors efficiently, it risks introducing unintended behavioral patterns, from rigid response styles to inappropriate social mimicry (e.g., expressing empathy or validating users in contexts where this can negatively influence performance and outcomes) [17, 80]. Thus, it may become difficult to selectively induce specific behaviors without introducing a broader set of human-like patterns.
@@
This approach impacts both interaction and evaluation. When interacting with models, users may develop anthropomorphic perceptions that lead to overreliance or emotional attachment, potentially interfering with goal-oriented tasks [3, 22]. During model evaluation, problematic feedback loops emerge when models trained with human-like traits are assessed using anthropomorphic signals. For instance, when evaluating if LLMs are “faking alignment," researchers might look for expressions of discomfort or hesitation, as a signal of misalignment [35, 6]. However, it is unclear if these signals genuinely reflect a model’s “internal state," or if they are merely learned behaviors resulting from post-training using human-like traits. This makes it challenging to distinguish “genuine” (mis)alignment from a learned performance of human-like discomfort or hesitation. Further research disaggregating the effects of post-training approaches on various evaluation outcomes can clarify and test whether these anthropomorphic signals provide meaningful information about model behavior, or if they primarily measure how well models have learned to simulate human-like responses.
@@
While current post-training techniques often default to human preferences as optimization targets, alternative frameworks could provide more precise specifications and compliance guarantees. Instead of aiming for human-like moral reasoning, we could focus on developing detailed, normative specifications, for example, based on the different roles (e.g., assistant vs teacher) AI systems play [91]. Instead of the anthropomorphic approach of instruction-tuning, recent work has demonstrated that non-anthropomorphic approaches (that do not include the step of providing an imperative “instruction” to the system as if speaking to a person) work as well for achieving model behavior on various tasks [38]. Control systems theory offers tools for maintaining system outputs within specified bounds, treating beneficial behavior as a problem of robust compliance rather than value alignment [10]. This becomes particularly crucial as models move beyond two-party interactions to more complex scenarios with multiple actors and potential adversarial inputs [73]. Advances in mechanistic interpretability techniques may also enable robust and direct verification and steering of model behavior against these specifications [13].
@@
4.3 Evaluation assumption: Model capabilities should be measured in human-like ways.
@@
As LLM developers have made rapid performance improvements, as assessed using various benchmarks, scholars have pointed out that current benchmarks can lead to incomplete or misguided understanding of model capabilities.
@@
Current LLM evaluations prioritize “black-box" behavioral testing analogous to the behaviorist paradigm in human psychology which measures performance primarily in the form of observable behaviors as opposed to mechanistic interpretation [19, 25]. Recent calls for a “science of evaluation" have formalized limitations in this approach, highlighting how current metrics and designs fall short of accounting for prompt sensitivity (e.g., dialect differences, punctuation and other small perturbations), the response structure of an evaluation (e.g., MCQ or open-ended response), generalization beyond a given test, as well as replicability [39]. Further such research quantifying the methodological limitations and error bounds of such evaluations can strengthen this behaviorist approach to measuring model capabilities [62]. Unlike humans, models can also quickly optimize for, and saturate, behavioral benchmarks without corresponding improvements in general capabilities. Yet, despite this pattern, many benchmarks remain static rather than being regularly refreshed, limiting their utility for meaningful evaluation [71]. Some recent work that challenges this assumption include efforts in dynamic benchmarking [50] and measuring performance in real-world LLM use contexts such as user-AI interactions [56, 18] to more accurately reflect model capabilities.
@@
Current evaluation frameworks predominantly rely on human performance benchmarks, from standardized tests (e.g., MMLU, GSM8K) to domain-specific examinations, as primary metrics for model capability assessment. This paradigm remains the main way progress is measured and communicated [76]. However, evaluating LLMs solely through human-centric tests risks overlooking LLMs unique strengths and weaknesses. McCoy et al. [59] argue that many current benchmarks drawn from tests designed to assess human cognition may highlight the overlap between human abilities and LLM capabilities while missing crucial failure modes specific to LLMs. They find robust evidence of failure modes in SOTA LLMs (including recent reasoning models like OpenAI’s o1) related to probabilities of examples and tasks [60]. This is because LLMs, trained on next-word prediction using massive text data, develop tendencies and biases that stem from their probabilistic training process. Drawing from cognitive science, they propose a “teleological approach" which “characterizes the problem that the system solves and to then use this characterization as a source of hypotheses about the system’s capacities and biases." In this case, given the problem is next-token prediction, they recommend designing tests which take into account sensitivities to task frequency in the training data as well as wording in prompts, among other things, to improves the predictive power of current LLM evaluation approaches [62].
4.4 Understanding model behavior assumption: Human-like normative judgments or intentions should be assigned to human-like model behaviors.
This assumption influences how we make sense of model behavior, particularly in how we assign fault, intention, and normative judgments to (i.e., consider good or bad) observed behaviors. The impact is especially notable in our understanding of failure modes. While models may exhibit seemingly human-like failure modes like sycophancy and hallucinations, framing these behaviors through human psychological concepts may constrain our solution space, by, for example, encouraging interventions that similarly rely on human psychological constructs (e.g., attempting to address sycophancy through prompts about independence or self-assertion).
@@
Hallucination is typically characterized as the problem of LLMs outputting factually incorrect information in a manner that suggests that they are true. Yet, this term obscures the mechanisms behind these phenomena: at risk of oversimplification, this behavior arises from the nature of language models as next-token predictors. Generated outputs are then labeled as hallucinations upon the reader’s normative judgment of whether or not they are useful, and not based on whether they are correct. Additionally, as Sui et al. [83] argue and show, what we commonly conceive of as hallucinations can actually be deeply valuable, and should not necessarily be dismissed as low-quality. They assert that hallucinations – or “confabulations”–should not be viewed as errors, but rather as particular model phenomena that offer unique benefits for applications like creativity, such as increased levels of narrativity [83, 28]. Yao et al. [89] also highlight that hallucinations ought to be viewed and utilized as adversarial examples rather than merely as bugs.
@@
The notion of sycophancy (i.e., the phenomena of LLM outputs that respond to the user’s input in ways that are perceived as overly servile, obedient, and/or flattering) [74, 80] is another example that reflects this assumption. Deciding whether an output is sycophantic or not is similarly a normative question: an output is sycophantic when it relates too closely to the prompt in ways that do not achieve the prompter’s goal. In contrast, recent work highlights how this property can be viewed as a strength: Li et al. [54] develop a methodology to use this mirroring to elicit, structure, and clarify users’ thinking across various task domains.
@@
The emerging body of research on LLM deception increasingly focuses on measuring strategic deception - defined as models “deceiving selectively based on incentives or instructions" [44]. While studies demonstrate that LLMs can produce deceptive statements in response to specific prompts, this work often faces two key interpretive challenges. First, it risks attributing observed behaviors to model intentions to deceive. Second, results are often interpreted as evidence of model-level deceptive traits rather than instance- and context-specific behaviors.
An alternative, less anthropomorphic framing, proposed by Shanahan et al. [79], views these behaviors through the lens of “role-play" where LLMs simulate human-like responses. This interpretation sees the system not as a singular entity but as context-bound, “inferring and applying approximate communicative intentions" [4]. Through this lens, complex behaviors like deception and self-awareness can be understood as sophisticated simulations rather than true cognitive states. This reframing also expands the set of interventions for deceptive behaviors: analyzing training data composition, examining how post-training interventions shape model behavior, and investigating reinforcement learning’s effects on output distributions.
4.5 User interaction assumption: Human-LLM interactions mirror human-human interactions.
@@
The de facto interaction paradigm for human-LLM interaction is prompt-based interfaces, originally designed as debugging tools for machine learning engineers [65]. As these interfaces resemble human-human chat interfaces, they may encourage users to naturally default to conversational patterns from human interaction. On one hand, this can reduce cognitive load and improve usability by drawing on user familiarity and acceptance of human-human chat interfaces [77]. However, research on effective prompting suggests that optimal results often require structured, sometimes non-intuitive formats (e.g., “least-to-most prompting") rather than human-like communication patterns which rely on shared context and paralinguistic cues [65, 92]. Simultaneously, research on human-LLM interaction shows that one of the key challenges users face is a significant gulf of envisioning or “distance between the human’s initial intentions and their formulation of a prompt that foresees how LLM capabilities and training data can be leveraged to generate high-quality output" [82]. This mismatch between natural dialogue and effective prompting requires greater experimentation with interaction paradigms and interface designs for LLMs. Structured interaction frameworks, using suggested inputs, guided flows, and/or domain-specific prompting strategies, would explicitly expose system capabilities rather than masking them behind conversational abstractions [82, 31, 30]. This can bridge the gulf of envisioning by enabling more systematic exploration of model functionality.